AIDA
GELINA
BRIKEN
nToF
CRIB
ISOLDE
CIRCE
nTOFCapture
DESPEC
DTAS
EDI_PSA
179Ta
CARME
StellarModelling
DCF
K40
CARME
Draft saved at 00:00:00
Fields marked with
*
are required
Entry time:
Sat Apr 5 01:28:14 2025
Author
*
:
Subject
*
:
<p>22:27</p> <p>Run 33 started cooled beam on target.</p> <p>Detectors IN (left arm pot =36.5 mm, right arm pot = 70.49 mm)</p> <p>Electron cooler voltage 754.945 this corresponds to 1.3499 MeV/u. This is above the resonance of interest but cooler settings cannot be altered without major beam tuning.</p> <p>Target density is 4.1e11 at/cm2, ring pressure looks OK YR09 1.2e-10 mbar.</p> <p>Horizontal beam position is -7.5 mm, vertical beam position is now 0.0 mm. This results in no beam halo in bottom left detector which should be good for seeing Rutherford.</p> <p>Beam is in the ring for 8 seconds and total cycle time between injections is 12 seconds</p> <p>FEE temperatures are OK. 3 FEEs are slightly too. (attachment 1)</p> <p>System checks are OK.</p> <p>Statistics look OK. High in the top left detector which is expected as majority of halo is there. (attachment 2+3)</p> <p>Merger OK Tape server OK - 851 kB/s</p> <p>Merger statistics (attachment 4)</p> <p>Leakage current is increasing but is likely temperature related. ML temp plot has same pattern and it is a WARM day. (attachment 5)</p> <p>Monitoring code - xy plot and Ey energy histograms (attachment 6+7). Rutherford rate looks higher -> looking peaky</p> <p>00:14</p> <p>System checks all OK</p> <p>FEE temperatures unchanged. Tom has now set up plotting of aida09 fee temperatures in the second workspace.</p> <p>Merger OK, tape server OK.</p> <p>Target density remains about 4.1e11 at/cm2. Ring pressure ~1e-10 mbar. YR08 read close to 1e-9 but beam is able to circulate in the ring so will not change for now.</p> <p>Rate statistics in all FEEs remain unchanged.</p> <p>02:10</p> <p>System checks OK</p> <p>FEE temperatures unchanged</p> <p>Merger OK</p> <p>Target density OK. Ring pressure not great but cannot change remotely due to script not in screen and poor coding.</p> <p>Beam stil on.</p> <p>Leakage current trending up.</p> <p> </p> <p>04.00 FEE64 temperatures - attachments 10-11</p> <p> aida09 ( and other FEE64s ) on downward trend</p> <p> CRYRING hall temperature - attachment 12</p> <p> </p> <p>05:00</p> <p>Ring pressure worsening and target getting thinner (4E11). Tweak nozzle when in the lab physically.</p> <p>Beam still on around 3E7</p> <p> </p> <p>Temperatures & Leakage currents OK. Same trends</p> <p>Merger continues OK</p> <p>Stats OK</p> <p>System wide checks OK</p> <p> </p> <p>07:00</p> <p>Fee temperatures have reduced - all green</p> <p>System check OK</p> <p>ADC data items statistics unchanged. </p> <p>Merger and tape server OK</p> <p>Leakage current plot (attachment 15)<br /> monitoring plots - xy plot and EY energy histogram (attachment 16+17 )</p> <p>Pressure has dropped slightly over the last hour, target density remains stable 4.2e11. Beam transformer has dropped off in the last hour -> less beam target interactions -> lower pressure?</p> <p>We still see count rates in the detectors.</p> <p>08:18</p> <p>Inconsistent beam into the ring continues. HKR are aware.</p> <p>Temperatures OK</p> <p>System checks OK</p> <p>10:00</p> <p>Beam is much more consistent now. We observe vast majority of shots injected into the ring and corresponding rate increase in the detectors via the monitoring code.</p> <p>FEE temperatures OK</p> <p>System wide checks OK</p> <p>Merger and tape server OK</p> <p>New merger stats (attachment 18)</p> <p>ADC data item statistics OK</p> <p>Leakage current plot (attachment 19)</p> <p>Attachments 20+21: Cooler settings before and after changing to the new beam energy of 1.3435 MeV/u (copy from GSI olog)</p> <p>10:24</p> <p>Run33 stopped. Run 34 started. Changing beam energy to be on resonance.</p> <p>Detectors moved fully out. Target moved out</p> <p>Final spectrum browser stat for all FEEs before run ended (attachment 22)</p>
Encoding
:
HTML
ELCode
plain
Suppress Email notification
Attachment 1:
Drop attachments here...
Draft saved at 00:00:00
ELOG V3.1.4-unknown